[2] Financial ratios (58 points) Find the missing values on the balance sheet and the income statement of Pensacola Railroads (in S millions). Start with the balance sheet, and then continue with the income statement. You can use the following ratios: NOTE: USE AVERAGE RECEIVABLES AND INVENTORY Debt Ratio loterest Coverage Ratio CurreatRatio Quick Ratio Cash Ratio 0.4 1.4 0.2 Average Collection Period (average receivables) Tax rate 73 das 40% Marks: 28 for balance sheet, 30 for income statement.Balance Sheet Pensacola Railroad 2017 2016 20 34 26 80 Cash & lnventory Current Assets Eixed Assets 25 Total assets 105 Notes Pavable. Accounts Pavable Current Liabilities, 25 20 30 35 20 Long termDebt Equity 30 115 105 Uncome 2017 Sales COGS Otber costs. 10 20 EBIT lnterest, Earnings before tax Tax. Net ocome[2] Financial ratios (58 points) NOTE: use average inventory and receivables Balance Sheet Pensacola Railroad Marks 2017 2016 Cash & marketable securities Accounts Receivable Inventory 20 34 26 80 Current Assets 25 105 Fixed Assets Total assets Notes Payable Accounts Payable 25 30 20 35 Current Liabilities Long term Debt Equity 20 30 Total Liabilities+Equity 115 105 28Income Statement 2017 Marks Sales COGS Other costs Depreciation EBIT Interest Earnings before tax Tax Net Income 10 20 30 Selected ratios Debt Ratio Interest Coverage Ratio Current Ratio Quick Ratio Cash Ratio Inventory Tumover (average inventory) Average Collection Period (average receivables) Tax rate 0.4 0.2 73 days 40%
Answer #1
2017 2016
Cash and marketable securities (55*0.2) 11.00 20.00
Accounts receivable = 55*1-11 = 44.00 34.00
Inventory 22.00 26.00
Current assets (55*1.4) 77.00 80.00
Fixed assets (115-77) 38.00 25.00
Total assets 115.00 105.00
Notes payable 25.00 20.00
Accounts payable 30.00 35.00
Current liabilities 55.00 55.00
Long term debt 6.38 20.00
Equity = (30.00+23.63) 53.63 30.00
Total liabilities+Equity 115.00 105.00
Sales 195.00
COGS 120.00
Other costs 10.00
Depreciation 20.00
EBIT 45.00
Interest (45/8) 5.63
Earnings before tax 39.38
Tax 15.75
Net income 23.63
WORKING NOTES FOR INCOME STATEMENT FIGURES:
Sales = Average receivables*365/73 = 39*365/73 = 195
COGS = Average inventory*5 = 24*5 = 120
Total assets = Total liabilities+Equity = 115
which is given.
Next find current assets = Current liabilities*Current ratio = 55*1.4 = 77
Further, cash = Current liabilities*0.2 = 55*0.2 = 11
Quick assets = Current liabilities*1 = 55
Accounts receivable = 55-cash = 55-11 = 44
Inventory = 77-11-44 = 22
Sales and COGS calculations are given.
Equity = beginning balance+Net income = 30+5.63
Long term debt = 115-53.63-55